OpenStudy (anonymous):

Find the sum of the finite geometric series. 3-9+27-81+243-729+2187-6561+19683-59049

OpenStudy (anonymous):

for the finite G series the formula is [a(1-r^(n-1)]/1-r

OpenStudy (anonymous):

a=3, r=-3 , n= 10 S(n) = a(1-r^(n) ) / (1-r) sub etc lol

OpenStudy (anonymous):

\[\frac {a(1-r^n) }{1-r}\]

OpenStudy (anonymous):

:( again mistake

OpenStudy (anonymous):

yes right :)

OpenStudy (anonymous):

wait min

OpenStudy (anonymous):

you might be right, yeh, thought it looked a little strange I wikipedia it, and it is a(1-r^(n-1) ) / (1-r)

OpenStudy (anonymous):

oh let me check

OpenStudy (dumbcow):

hmm no you were right before

OpenStudy (anonymous):

no I was correct firsdt time

OpenStudy (anonymous):

yes , just went to check a high school maths study guide lying around

OpenStudy (anonymous):

yes the first 1 is right

OpenStudy (anonymous):

damm wikipedia, its usually a good resource for maths

OpenStudy (anonymous):

but wikipedia shows the same formula u wrote at the first place :)

OpenStudy (anonymous):

I can seem to find it, ahh I am bit tired

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

i am lost now lol

OpenStudy (anonymous):

yes its for sure :)

OpenStudy (anonymous):

\[\frac {a(1-r^n)}{1-r}\] sub in a=3, r=-3, n=10

OpenStudy (anonymous):

3[1-(-3)^10]/1-3 =3[1-3^10]/2

OpenStudy (anonymous):

so is the answer 1536?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

u calculated it?does it come 1536?

OpenStudy (anonymous):

yes that is what i got, is that what you got?

OpenStudy (anonymous):

i dint calculate :)

OpenStudy (anonymous):

oh ok... but does that make sense? my answer???

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

lol ok thank you